# Improved QLDPC Surgery: Logical Measurements and Bridging Codes

### v3

@@ 1 Introduction
Recently, several protocols for quantum memories based on LDPC codes have been developed [BVC+17, PNFB17, HB23, BCG+24, XBAP+24]. Efficient memories could be used in concert with the surface code to reduce space overhead. In some instances, these memory protocols simultaneously have high thresholds and low overheads at practical code block sizes, making them attractive candidates for near-term demonstrations [BCG+24]. A growing body of research aims to find improved codes [SHR24, VXHB24], decoders [GCR24, WB24], and implementations for these quantum memories [VYL+23, PGP+24, BDS+24].
@@
In an earlier work [CKBB22], Cohen, Kim, Bartlett, and Brown introduced a scheme based on an ancilla system which is a hypergraph product code. We henceforth refer to their scheme as the CKBB scheme and explain its mechanism in Section 2.2. For a logical operator of weight $d$, the CKBB ancilla system has $12d-$ 1 layers, each layer isomorphic to the Tanner graph of the measured operator, totaling to a space overhead of $O(d^{2})$. This ancilla system can be naturally extended to support joint measurements. The scheme then use $d$ rounds of QEC on the merged code to obtain the logical measurement result. If the original code has qubit degree $\delta$ and distance $d$, the merged code will have degree $\delta+1$ and distance $d$. While the CKBB scheme has many desirable properties, its space overhead is daunting: for typical LDPC codes where the distance scales as $d\sim\Omega(\sqrt{n})$, the ancilla system needed for measuring one logical operator may be larger than the code block itself. This poses a significant barrier to utilizing the CKBB scheme in practice.
@@
    We gauge-fix the CKBB ancilla system so that the merged code contains no gauge qubits. The resulting gauge-fixed ancilla system ensures that the merged code has distance at least $d$, while using much fewer than $12d-$ 1 layers if the Tanner graph induced by the logical operator is expanding (Theorem 6). In cases where the graph is sufficiently expanding, the space overhead for measuring a weight $w$ logical operator is just $\Theta(w)$.
@@
Our gauge-fixed surgery scheme comes with a noteworthy caveat: without additional assumptions on the measured operator or the code, we do not have upper bounds on the weight of the gauge-fixed operators. In other words, gauge-fixing these operators may break the LDPC property on the merged code. We discuss this caveat at length in Section 3.4, where we show theoretical evidence (Lemma 12) suggesting that for many important families of QLDPC codes, these gauge checks may not break the LDPC property. Notably, we prove that for hypergraph product codes the merged code has degree $\delta+1$ if the original code has degree $\delta$ (Theorem 13). Our bridge construction also lacks a worst case guarantee on the LDPC property, which we discuss in Section 3.8. We remark that after this work first appeared, these caveats are fully resolved in the schemes proposed by [WY24] and [SJOY24]. See section 1.1 for a more detailed discussion on related works.
@@ 1.1 Related work
Since this work first appeared, a few others have followed with refinements to the central ideas. First, while we point out here that expansion existing in the Tanner graph of the code can drastically reduce the space-overhead of QLDPC surgery, the techniques of Refs. [WY24] and [IGND24] can simply add the required expansion to the ancilla system to achieve the same or better space overhead. In addition, Refs. [WY24, IGND24] think of the ancilla system as built on a graph and this perspective has other benefits in that certain cellulation [Has21, WY24, IGND24] and decongestion [FH21, Has21, WY24] ideas can be applied to guarantee that the gauge-fixed ancilla system remains LDPC. Second, while we point out here that any two ancilla systems can be joined via an adapter to measure a joint logical operator, Ref. [SJOY24] constructs the adapter in such a way to guarantee the newly introduced checks are also still LDPC. This construction relies also on the graphical perspective from Refs. [WY24, IGND24]. Third, both Refs. [XZZ+24] and [ZL24] present approaches to parallelizing related logical operations and appeared around the same time as our work. Ref. [XZZ+24] does this specifically for homological product codes using the similar technique of homomorphic measurements [HJOY23], while Ref. [ZL24] creates some parallelized surgery methods in the CKBB [CKBB22] style without gauge-fixing.
The independent work of [Cow24] appeared shortly before ours, in which QLDPC surgery is studied in the language of chain complexes and homology. [Cow24] further applied the CKBB construction to various small-scale QLDPC codes, and presented numerical results showing that in concrete instances, the space-overhead is often much less than the $O(d^{2})$ upper bound, which is reflected here in Table 1.
@@ 2.1 Tanner graph notation
We describe quantum codes with or without ancilla systems by drawing their Tanner graphs. Instead of drawing the Tanner graphs in their entirety, we combine nodes into several sets of either check nodes or qubit nodes. These sets should be chosen so that we can easily describe the connections between each pair of qubit and check sets via an incidence matrix. To be more specific, if there is a set of qubits $V$ and a set of checks $C$, the edge between them will be labeled by a matrix $F\in\mathbb{F}_{2}^{|C|\times|V|}$ where $F_{ij}=1$ if and only if check $i$ from $C$ is connected to qubit $j$ from $V$. If these checks are $Z$ checks, then we use $\to_{Z}$ to summarize the above situation as $F:C\to_{Z}V$. For example the code $\mathcal{G}$ above has connections $H^{X}:\mathcal{C}^{X}\to_{X}\mathcal{V}$ and $H^{Z}:\mathcal{C}^{Z}\to_{Z}\mathcal{V}$. Since our graphs are sparse, many pairs of qubit and check sets will not be connected (or, formally, connected by the 00 matrix), in which case we leave out that edge.
@@ 2.2 The CKBB ancilla system
A pair of $V_{i},C_{i}$, connected as in the Tanner graph constitute one ‘layer’ of the construction. Having identified $C_{0},V_{0}$ within the original code $\mathcal{G}$’s vertices $\mathcal{C}^{Z},\mathcal{V}$, we append $12d-$ 1 more such layers. We assign these vertices alternating roles, as follows.
$V_{i}\text{ for }\Bigg{\{}\begin{array}[]{l}\text{odd }i\to\text{ X check}\\ \text{even }i\to\text{ physical qubit}\end{array}$ (3)
$C_{i}\text{ for }\Bigg{\{}\begin{array}[]{l}\text{odd }i\to\text{ physical qubit}\\ \text{even }i\to\text{ Z check}\end{array}$ (6)
Layers are connected identically via $F:C_{i}\to_{Z}V_{i}$ for even $i$, and $F^{\top}:V_{i}\to_{X}C_{i}$ for odd $i$. Furthermore, adjacent layers are connected by the identity: $I:C_{0}\to_{Z}C_{1}$, $I:V_{1}\to_{X}C_{0}$, $I:V_{1}\to_{X}C_{2}$, $I:C_{2}\to_{Z}C_{1}$, etc. This construction is visualized in Figure 1.
@@
In the resulting code the operator $\bar{X}_{M}$ is a product of stabilizers, namely all the appended $X$-checks $V_{1},V_{3},...,V_{2d-1}$. Therefore a code-switching scheme from $\mathcal{G}$ to the larger code can be used for a fault-tolerant logical measurement of $\bar{X}_{M}$. If any other logical operators were contained in the support of $\bar{X}_{M}$, then these would be measured as well. Therefore it is key that $\bar{X}_{M}$ is irreducible.
@@ 3.1 Gauge-fixed $X$ Ancilla Systems
Attaching such an ancilla system introduces new stabilizers and gauge degrees of freedom to the code, which could reduce the code distance. Specifically, multiplying a logical operator of the original code by new stabilizers or gauge operators could reduce its weight below $d$. In [CKBB22], this issue is resolved by taking $L=2d-1$.
@@
We defer the proof of this theorem to Appendix A.1. We note that without additional knowledge on the structure of the code or the measured operator, we do not have upper bounds on the row or column weight of $G$. While this is an undesirable caveat, we show theoretical evidence in Section 3.4 suggesting that for many important families of QLDPC codes, these gauge checks may not break the LDPC property. In particular, the gauge checks may be products of other $Z$ stabilizers of $\mathcal{G}_{X}$ and hence do not need to be explicitly measured. This happens for the gross code as we will see in Section 4 and for hypergraph product codes, as we will show in Theorem 13.
@@ 3.2 $X$ Measurement Protocol
$\mathcal{D}$ splits into steps $\mathcal{D}_{0},...,\mathcal{D}_{R}$, corresponding to $R$ measurements of the checks of $\mathcal{G}_{X}$. The first and final steps $\mathcal{D}_{0}$ and $\mathcal{D}_{d}$ handle the code deformation between $\mathcal{G}$ and $\mathcal{G}_{X}$. As shown in Figure 2, we call the first layer ($V_{1},C_{1}$) of $\mathcal{G}_{X}$ the interface, and the remaining layers are the module. There is no module when $L=1$, so the gauge checks $U_{1}$ are considered part of the interface. We call a check unreliable if, even in the noise-free case, the outcome of the check is random since it doesn’t commute with the stabilizer of the current state. A reliable check is deterministic in the noise-free case, but not necessarily $+1$.
@@
    Initialization: Initialize the interface qubits, namely qubits in $C_{1}$, to |$\ket{0}$. If there is a module ($L>1$), initialize a stabilizer state on the module qubits $C_{2},V_{3},\cdots,V_{L}$. It has the stabilizers $H_{\text{mod}}^{Z}(C_{2})$, $\mathcal{H}^{Z}(C_{i})$ for even $i\geq 4$, $\mathcal{H}^{Z}(U_{L})$, and $\mathcal{H}^{X}(V_{i})$ for odd $i\geq 3$. We refer to these checks as the module stabilizer.
    Measure the module stabilizers. Additionally, either measure the stabilizers of $\mathcal{G}$, or take the measurement outcomes from a prior round. If there is no module ($L=1$), then we can still compute gauge checks $\mathcal{H}^{Z}(U_{1})$ by measuring the interface qubits in the $Z$ basis and computing the appropriate parities.
@@
    Fusion: Measure the stabilizers of $\mathcal{G}_{X}$. Of these, the $\mathcal{H}^{X}(V_{1})$ checks are unreliable, but everything else is reliable. Let $\mathcal{D}_{0}$ contain parities between pairs of reliable checks from step 1 and step 2 only. This involves taking parities between step 1’s $H^{Z}(C_{0})$ and step 2’s $\mathcal{H}^{Z}(C_{0})$, as well as step 1’s $H_{\text{mod}}^{Z}(C_{2})$ and step 2’s $\mathcal{H}^{Z}(C_{2})$. The module stabilizers except $H_{\text{mod}}^{Z}(C_{2})$ are identical in step 1 and step 2 so these parities are zero. From round 1 to 2, checks $H^{Z}(C_{0})$ and $H_{\text{mod}}^{Z}(C_{2})$ have gained support on the interface $C_{1}$ and become $\mathcal{H}^{Z}(C_{0})$ and $\mathcal{H}^{Z}(C_{2})$. But since the interface qubits are in the |$\ket{0}$ state, the measurement outcome is unaffected, so these parities should be zero as well.
@@ 3.3 Code Distance and Fault Distance
Any $X$-type logical operator of code $\mathcal{G}$ is still a logical operator of the merged code $\mathcal{G}_{X}$. Irreducibility of $\bar{X}_{M}$ implies that this set of logical $X$-type operators provides a complete basis of $X$-type logical operators of the $k-1$ qubit code $\mathcal{G}_{X}$. Hence, consider any nontrivial $X$-type logical operator $P$ of $\mathcal{G}_{X}$ supported entirely on $\mathcal{V}$. It is also a logical operator of $\mathcal{G}$ so it has weight at least $d$. We will show $P$ cannot be reduced in weight to below $d$ by multiplying by the $X$-checks of $\mathcal{G}_{X}$.
@@
Let $s=t\cap|V_{0}|$, from Lemma 4, we see that $|P^{\prime}|\geq|t|-|s|+\min(|s|,|V_{0}|-|s|)$. If $|s|\leq|V_{0}|/2$, we see that $|P^{\prime}|\geq|t|\geq d$. Otherwise, we have
@@
As a corollary, we see that the space overhead of our ancilla system is $O(d/\beta)$. We remark that boundary expansion is a demanding property, and for general codes we expect $1/\beta$ to be super-constant. However, in the case where every check in the induced Tanner graph acts on exactly two qubits in the support of $\bar{X}_{M}$, the induce Tanner graph can be interpreted as a simple graph by turning every check into an edge. Boundary Cheeger constant on the induced Tanner graph then translates to the usual Cheeger constant on this simple graph. Since edge expansion is a much more common property, for such codes the space overhead may be linear. We note that in [WY24] and [IGND24], instead of using the induced Tanner graph (which is generally a hypergraph if we treat each induced check as a hyperedge) for measurement, the proposed schemes constructed and used a compatible expander graph with sufficient edge expansion (Cheeger constant $\geq 1$) for measurement. This novelty guarantees a low space-overhead ancilla system for any QLDPC code. We also note that in [SJOY24], the requirement on edge expansion is further relaxed to a notion called relative expansion.
@@
Consider a phenomenological noise model on the detector graph $\mathcal{D}$: Pauli errors can occur on qubits in between any rounds of measurement, and any measurement outcome can be faulty. We say a detector triggers when the associated parity between measurement outcomes is 1, and we say an error is undetected if it triggers no detectors. The measurement fault distance is the minimum weight of an undetected error that causes an error in the measurement outcome. The logical fault distance is the minimum weight of an undetected error that causes a logical Pauli error on a qubit in $\mathcal{G}$ other than the measured qubit and does not cause a measurement error.
@@
and the stabilizer group of the original code, interface qubits in |$\ket{0}$, and module, with or without $\bar{X}_{M}$ included
@@
All three codes corresponding to $\mathcal{S}^{\prime}$, $\mathcal{S}_{X}$, and $\mathcal{U}$ have $k-1$ logical qubits and a logical basis for each can be chosen to consist of the same set of operators $\mathcal{L}=\langle\bar{X}_{i},\bar{Z}_{i}:i=1,\dots,k-1\rangle$. To see this, note that any logical $X$ operator $\bar{X}_{i}$ (inequivalent to $\bar{X}_{M}$) of $\mathcal{S}$ on the original code qubits is still a logical operator of each of the three codes. Also, if $\bar{Z}^{\prime}_{i}$ is a nontrivial logical $Z$ operator of $\mathcal{S}$ that commutes with $\bar{X}_{M}$, then we claim it can be multiplied by stabilizers of the original code to obtain an equivalent $\bar{Z}_{i}$ suitable for inclusion in $\mathcal{L}$ because it has no support on $V_{0}$. This statement is a corollary of the following lemma.
@@
We also use =$\mathcal{L}^{*}=\mathcal{L}\setminus\{I\}$ to denote the nontrivial logical operators.
@@
We denote by $\mathcal{M}_{t}$ the (abelian) group of Paulis generated by the checks measured in timestep $t$. In particular, $\mathcal{M}_{0}=\mathcal{S}=\mathcal{M}_{R+1}$ measures the original code, interface qubits in |$\ket{0}$, and module (see the top and bottom of Figure 3). For $10<t<R+$ 1, $\mathcal{M}_{t}$ measures the merged code stabilizers $\mathcal{S}_{X}$.
@@
Detector groups $\mathcal{M}_{t\rightarrow t^{\prime}}$ fall into three cases. If $t=0$ and $10<t^{\prime}<R+$ 1 or $10<t<R+$ 1 and $t^{\prime}=R+1$, then $\mathcal{M}_{t\rightarrow t^{\prime}}=\mathcal{S}\cap\mathcal{S}_{X}$, the stabilizers that are both in the original code and merged code. If $10<t,t^{\prime}<R+$ 1, then $\mathcal{M}_{t\rightarrow t^{\prime}}=\mathcal{S}_{X}$, the merged code stabilizers. Finally, if $t=0$ and $t^{\prime}=R+1$, then $\mathcal{M}_{0\rightarrow R+1}=\mathcal{S}$, the original code stabilizers.
We assume that the final round is free of measurement errors to disallow hiding detectors at the future time boundary. Presumably this strategy would be unsuccessful anyway since it would be detected by future $t>R+1$ measurements. Because the final round is perfect, any qubit errors must at this point amount to an undetectable error. We point out that an ideal round of measurement is also done in our numerical case study (see Section 4) at $t=R+2$ for the same reason.
@@
The elements highlighted in gray are a generating set of $\mathcal{U}$ and the $\bar{X}_{i}$ and $\bar{Z}_{i}$ in the last column are the generators of $\mathcal{L}$. A measurement fault is an element of $\mathcal{C}(\mathcal{S}\cap\mathcal{S}_{X})$ that includes $\bar{Z}_{M}$, the only generator of $\mathcal{C}(\mathcal{S}\cap\mathcal{S}_{X})$ that anticommutes with $\bar{X}_{M}$. So measurement faults are exactly the elements of $\bar{Z}_{M}\mathcal{L}\mathcal{U}$. Similarly, logical faults are elements of $\mathcal{C}(\mathcal{S}\cap\mathcal{S}_{X})$ that do not include $\bar{Z}_{M}$ (since Definition 7 specifies that logical faults do not flip the logical measurement) and include at least one of $\bar{X}_{i}$ or $\bar{Z}_{i}$. Thus, logical faults are exactly the elements of $\mathcal{L}^{*}\mathcal{U}$.
@@
The highlighted elements are generators of $\mathcal{S}^{\prime}$. Similar to above, measurement faults are exactly elements of $\bar{Z}_{M}\mathcal{L}\mathcal{S}^{\prime}$ and logical faults are exactly elements of $\mathcal{L}^{*}\mathcal{S}^{\prime}$.

### v4

@@ 1 Introduction
Recently, several protocols for quantum memories based on LDPC codes have been developed [BVC+17, PNFB17, HB24, BCG+24, XBAP+24]. Efficient memories could be used in concert with the surface code to reduce space overhead. In some instances, these memory protocols simultaneously have high thresholds and low overheads at practical code block sizes, making them attractive candidates for near-term demonstrations [BCG+24]. A growing body of research aims to find improved codes [SHR24, VXHB25], decoders [GCR24, WB24], and implementations for these quantum memories [VYL+23, PGP+25, BDS+25].
@@
In an earlier work [CKBB22], Cohen, Kim, Bartlett, and Brown introduced a scheme based on an ancilla system which is a hypergraph product code. We henceforth refer to their scheme as the CKBB scheme and explain its mechanism in Section 2.2. For a logical operator of weight $d$, the CKBB ancilla system has 2d12d-1 layers, each layer isomorphic to the Tanner graph of the measured operator, totaling to a space overhead of $O(d^{2})$. This ancilla system can be naturally extended to support joint measurements. The scheme then use $d$ rounds of QEC on the merged code to obtain the logical measurement result. If the original code has qubit degree $\delta$ and distance $d$, the merged code will have degree $\delta+1$ and distance $d$. While the CKBB scheme has many desirable properties, its space overhead is daunting: for typical LDPC codes where the distance scales as $d\sim\Omega(\sqrt{n})$, the ancilla system needed for measuring one logical operator may be larger than the code block itself. This poses a significant barrier to utilizing the CKBB scheme in practice.
@@
    We gauge-fix the CKBB ancilla system so that the merged code contains no gauge qubits. The resulting gauge-fixed ancilla system ensures that the merged code has distance at least $d$, while using much fewer than 2d12d-1 layers if the Tanner graph induced by the logical operator is expanding (Theorem 6). In cases where the graph is sufficiently expanding, the space overhead for measuring a weight $w$ logical operator is just $\Theta(w)$.
@@
Our gauge-fixed surgery scheme comes with a noteworthy caveat: without additional assumptions on the measured operator or the code, we do not have upper bounds on the weight of the gauge-fixed operators. In other words, gauge-fixing these operators may break the LDPC property on the merged code. We discuss this caveat in Section 3.4, where we show theoretical evidence (Lemma 12) suggesting that in the average case, these gauge checks may not break the LDPC property. Our bridge construction also lacks a worst case guarantee on the LDPC property, which we discuss in Section 3.8. We remark that after this work first appeared, these caveats are fully resolved in the schemes proposed by [WY24] and [SJOY24]. See section 1.1 for a more detailed discussion on related works.
@@ 1.1 Related work
We remark on a collection of related works. The independent work of [Cow24] appeared shortly before ours, in which QLDPC surgery is studied in the language of chain complexes and homology. [Cow24] further applied the CKBB construction to various small-scale QLDPC codes, and presented numerical results showing that in concrete instances, the space-overhead is often much less than the $O(d^{2})$ upper bound, which is reflected here in Table 1.
Since this work first appeared, a few others have followed with improvements to the central ideas. First, while we point out here that expansion existing in the Tanner graph of the code can drastically reduce the space-overhead of QLDPC surgery, the techniques of [WY24] and [IGND25] can simply add the required expansion to the ancilla system to achieve the same or better space overhead. In addition, [WY24, IGND25] think of the ancilla system as built on a graph and this perspective has other benefits in that certain cellulation [Has21, WY24, IGND25] and decongestion [FH21, Has21, WY24] ideas can be applied to guarantee that the gauge-fixed ancilla system remains LDPC. Second, while we show that any two ancilla systems can be joined via a bridge to measure a joint logical operator, [SJOY24] constructs the bridge/adapter in such a way to guarantee the newly introduced checks are still LDPC. This construction relies also on the graphical perspective from [WY24, IGND25]. Third, both [XZZ+25] and [ZL25] present approaches to parallelizing related logical operations and appeared around the same time as our work. Ref. [XZZ+25] applied the technique of homomorphic measurements [HJOY23] to homological product codes, while Ref. [ZL25] proposed parallelized QLDPC surgery methods applicable to generic codes, utilizing the CKBB scheme. Building on the methods of our work and [WY24, SJOY24][HCWY25] and [YSR+25] proposed QLDPC architectures capable of performing universal fault-tolerant computation with fixed hardware connectivity, while [CHWY25] improved the techniques from [ZL25] to enable more efficient parallel surgery. We refer readers to Section 3.2 of [HCWY25] for a brief review of recent developments.
@@ 2.1 Tanner graph notation
We describe quantum codes with or without ancilla systems by drawing their Tanner graphs. Instead of drawing the Tanner graphs in their entirety, we combine nodes into several sets of either check nodes or qubit nodes. These sets should be chosen so that we can easily describe the connections between each pair of qubit and check sets via an incidence matrix. To be more specific, if there is a set of qubits $V$ and a set of checks $C$, the edge between them will be labeled by a matrix $F\in\mathbb{F}_{2}^{|C|\times|V|}$ where $F_{ij}=1$ if and only if check $i$ from $C$ is connected to qubit $j$ from $V$. If these checks are $Z$ checks, then we use $\to_{Z}$ to summarize the above situation as $F:C\to_{Z}V$. For example the code $\mathcal{G}$ above has connections $H^{X}:\mathcal{C}^{X}\to_{X}\mathcal{V}$ and $H^{Z}:\mathcal{C}^{Z}\to_{Z}\mathcal{V}$. Since our graphs are sparse, many pairs of qubit and check sets will not be connected (or, formally, connected by the 0 matrix), in which case we leave out that edge.
@@ 2.2 The CKBB ancilla system
A pair of $V_{i},C_{i}$, connected as in the Tanner graph constitute one ‘layer’ of the construction. Having identified $C_{0},V_{0}$ within the original code $\mathcal{G}$’s vertices $\mathcal{C}^{Z},\mathcal{V}$, we append 2d12d-1 more such layers. We assign these vertices alternating roles, as follows.
[…]
Layers are connected identically via $F:C_{i}\to_{Z}V_{i}$ for even $i$, and $F^{\top}:V_{i}\to_{X}C_{i}$ for odd $i$. Furthermore, adjacent layers are connected by the identity: $I:C_{0}\to_{Z}C_{1}$, $I:V_{1}\to_{X}V_{0}$, $I:V_{1}\to_{X}V_{2}$, $I:C_{2}\to_{Z}C_{1}$, etc. This construction is visualized in Figure 1.
@@
In the resulting code the operator $\bar{X}_{M}$ is a product of stabilizers, namely all the appended $X$-checks $V_{1},V_{3},...,V_{2d-1}$. Therefore a code-switching scheme from $\mathcal{G}$ to the larger code can be used for a fault-tolerant logical measurement of $\bar{X}_{M}$. If any other $X$ logical operators were contained in the support of $\bar{X}_{M}$, then these would be measured as well. Therefore it is key that $\bar{X}_{M}$ is irreducible.
@@ 3.1 Gauge-fixed $X$ Ancilla Systems
Attaching such an ancilla system introduces new stabilizers and gauge degrees of freedom to the code, which could reduce the code distance. Specifically, multiplying a logical operator of the original code by new stabilizers or gauge operators could reduce its weight below $d$. In [CKBB22], this issue is resolved by taking L=2d1L=2d-1.
@@
We defer the proof of this theorem to Appendix A.1. We note that without additional knowledge on the structure of the code or the measured operator, we do not have upper bounds on the row or column weight of $G$. While this is an undesirable caveat, we show theoretical evidence in Section 3.4 suggesting that in the average case, these gauge checks may not break the LDPC property. In particular, the gauge checks may be products of other $Z$ stabilizers of $\mathcal{G}_{X}$ and hence do not need to be explicitly measured.
@@ 3.2 $X$ Measurement Protocol
$\mathcal{D}$ splits into steps $\mathcal{D}_{0},...,\mathcal{D}_{R}$, corresponding to $R$ measurements of the checks of $\mathcal{G}_{X}$. The first and final steps $\mathcal{D}_{0}$ and $\mathcal{D}_{d}$ handle the code deformation between $\mathcal{G}$ and $\mathcal{G}_{X}$. As shown in Figure 2, we call the first layer ($V_{1},C_{1}$) of $\mathcal{G}_{X}$ the interface, and the remaining layers are the module. There is no module when $L=1$, so the gauge checks $U_{1}$ are considered part of the interface. We call a check unreliable if, even in the noise-free case, the outcome of the check is random since it doesn’t commute with the stabilizer of the current state. A reliable check is deterministic in the noise-free case, but not necessarily +1+1.
@@
[…]
    Measure the module stabilizers. Additionally, either measure the stabilizers of $\mathcal{G}$, or take the measurement outcomes from a prior round.
@@
[…]
@@ 3.3 Code Distance and Fault Distance
Any $X$-type logical operator of code $\mathcal{G}$ is still a logical operator of the merged code $\mathcal{G}_{X}$. Irreducibility of $\bar{X}_{M}$ implies that this set of logical $X$-type operators provides a complete basis of $X$-type logical operators of the k1k-1 qubit code $\mathcal{G}_{X}$. Hence, consider any nontrivial $X$-type logical operator $P$ of $\mathcal{G}_{X}$ supported entirely on $\mathcal{V}$. It is also a logical operator of $\mathcal{G}$ so it has weight at least $d$. We will show $P$ cannot be reduced in weight to below $d$ by multiplying by the $X$-checks of $\mathcal{G}_{X}$.
@@
Let $s=t\cap V_{0}$, from Lemma 4, we see that $|P^{\prime}|\geq|t|-|s|+\min(|s|,|V_{0}|-|s|)$. If $|s|\leq|V_{0}|/2$, we see that $|P^{\prime}|\geq|t|\geq d$. Otherwise, we have
@@
As a corollary, we see that the space overhead of our ancilla system is $O(d/\beta)$. We remark that boundary expansion is a demanding property, and for general codes we expect $1/\beta$ to be super-constant. However, in the case where every check in the induced Tanner graph acts on exactly two qubits in the support of $\bar{X}_{M}$, the induce Tanner graph can be interpreted as a simple graph by turning every check into an edge. Boundary Cheeger constant on the induced Tanner graph then translates to the usual Cheeger constant on this simple graph. Since edge expansion is a much more common property, for such codes the space overhead may be linear. We note that in [WY24] and [IGND25], instead of using the induced Tanner graph (which is generally a hypergraph if we treat each induced check as a hyperedge) for measurement, the proposed schemes constructed and used a compatible expander graph with sufficient edge expansion (Cheeger constant $\geq 1$) for measurement. This novelty guarantees a low space-overhead ancilla system for any QLDPC code. We also note that in [SJOY24], the requirement on edge expansion is further relaxed to a notion called relative expansion.
@@
Consider a phenomenological noise model on the detector graph $\mathcal{D}$: Pauli errors can occur on qubits in between any rounds of measurement, and any measurement outcome can be faulty. We say a detector triggers when the associated parity between measurement outcomes is 1, and we say an error is undetected if it triggers no detectors. The logical fault distance is the minimum weight of an undetected error that causes a logical Pauli error on a qubit in $\mathcal{G}$ other than the measured qubit and does not cause a measurement error. The measurement fault distance is the minimum weight of an undetected error that causes an error in the measurement outcome.
@@
[…]
@@
All three codes corresponding to $\mathcal{S}^{\prime}$, $\mathcal{S}_{X}$, and $\mathcal{U}$ have k1k-1 logical qubits and a logical basis for each can be chosen to consist of the same set of operators $\mathcal{L}=\langle\bar{X}_{i},\bar{Z}_{i}:i=1,\dots,k-1\rangle$. To see this, note that any logical $X$ operator $\bar{X}_{i}$ (inequivalent to $\bar{X}_{M}$) of $\mathcal{S}$ on the original code qubits is still a logical operator of each of the three codes. Also, if $\bar{Z}^{\prime}_{i}$ is a nontrivial logical $Z$ operator of $\mathcal{S}$ that commutes with $\bar{X}_{M}$, then we claim it can be multiplied by stabilizers of the original code to obtain an equivalent $\bar{Z}_{i}$ suitable for inclusion in $\mathcal{L}$ because it has no support on $V_{0}$. This statement is a corollary of the following lemma.
@@
We also use $\mathcal{L}^{*}=\mathcal{L}\setminus\{I\}$ to denote the nontrivial logical operators.
@@
[…]
@@
Detector groups $\mathcal{M}_{t\rightarrow t^{\prime}}$ fall into three cases. If $t=0$ and $0<t^{\prime}<R+1$ or $0<t<R+1$ and $t^{\prime}=R+1$, then $\mathcal{M}_{t\rightarrow t^{\prime}}=\mathcal{S}\cap\mathcal{S}_{X}$, the stabilizers that are both in the original code and merged code. If $0<t,t^{\prime}<R+1$, then $\mathcal{M}_{t\rightarrow t^{\prime}}=\mathcal{S}_{X}$, the merged code stabilizers. Finally, if $t=0$ and $t^{\prime}=R+1$, then $\mathcal{M}_{0\rightarrow R+1}=\mathcal{S}$, the original code stabilizers.
We make the standard modeling assumption that the final round is free of measurement errors, see Ref. [FSG09], for instance. Because the final round is perfect, any qubit errors must at this point amount to an undetectable error, either a stabilizer or non-trivial logical error, which makes the definition of a logical failure clear-cut. We point out that an ideal round of measurement is also done in our numerical case study (see Section 4) at $t=R+2$ for the same reason.
@@
The elements highlighted in gray are a generating set of $\mathcal{U}$ and the $\bar{X}_{i}$ and $\bar{Z}_{i}$ in the last column are the generators of $\mathcal{L}$. A measurement fault is an element of $\mathcal{C}(\mathcal{S}\cap\mathcal{S}_{X})$ that includes $\bar{Z}_{M}$, the only generator of $\mathcal{C}(\mathcal{S}\cap\mathcal{S}_{X})$ that anticommutes with $\bar{X}_{M}$. So, for this first case, measurement faults are exactly the elements of $\bar{Z}_{M}\mathcal{L}\mathcal{U}$. Similarly, logical faults are elements of $\mathcal{C}(\mathcal{S}\cap\mathcal{S}_{X})$ that do not include $\bar{Z}_{M}$ (since Definition 7 specifies that logical faults do not flip the logical measurement) and include at least one of $\bar{X}_{i}$ or $\bar{Z}_{i}$. Thus, logical faults are exactly the elements of $\mathcal{L}^{*}\mathcal{U}$.
@@
The highlighted elements are generators of $\mathcal{S}^{\prime}$. Similar to the first case above, measurement faults in this third case are exactly elements of $\bar{Z}_{M}\mathcal{L}\mathcal{S}^{\prime}$ and logical faults are exactly elements of $\mathcal{L}^{*}\mathcal{S}^{\prime}$.